OpenStudy (anonymous):

Write the equation with rectangular coordinates. r^2=6cos(2theta) (x2 + y2)2 =

zepdrix (zepdrix):

Hmmmm

zepdrix (zepdrix):

So you're trying to solve for \(\Large\rm (x^2+y^2)^2\) ? Weird :U Ok ok ok ok ok here is what I'm thinkin...

zepdrix (zepdrix):

\[\Large\rm r^2=6\cos(2\theta)\]Multiply both sides by r^2,\[\Large\rm (r^2)^2=6r^2 \cos(2\theta)\]Rewrite the right side using our Cosine Double Angle Identity:\[\Large\rm (r^2)^2=6r^2(\cos^2\theta-\sin^2\theta)\]Distribute the r^2,\[\Large\rm (r^2)^2=6r^2\cos^2\theta-6r^2\sin^2\theta\]

zepdrix (zepdrix):

Then use our things for converting from polar to rectangular:\[\Large\rm x=r \cos \theta, \qquad y=r \sin \theta\]

zepdrix (zepdrix):

Giving us,\[\Large\rm (x^2+y^2)^2=6x^2-6y^2\]Something like that I guess....?

OpenStudy (anonymous):

wow sorry i hadnt seen that someone had responded

OpenStudy (anonymous):

wow thank you